You place a​ stop-loss order to sell at $ 52.00 on a stock currently selling for $55.96 per​ share.

What is likely to be the minimum loss you will experience on 100 shares if the stock price rapidly declines to $49.17 per​ share? Round to the neared cent

What if you had placed a​ stop-limit order to sell at $52.00 and the stock price tumbled to $49.17​? Round to the nearest cent
Answer #1

As you have put a stop loss @$52, you will have minimum loss when your trade will get executed @$52 when price rapidly declines to $49.17.

Your minimum loss= 100*(55.96-52)=$396

If you had put a stop limit order to sell @$52, your trade will get executed at current market price i.e. $55.96

Here your loss=100*(55.96-55.96)=0